OpenStudy (anonymous):

The speed of sound in the air is proportional to the square root of the absolute temperature. At 20 °C the speed of sound is 343 m/s. What is the speed of sound at –10 °C? (0 °C = 273 K).

OpenStudy (anonymous):

please, help

OpenStudy (anonymous):

v=343*sqrt(263/293)

OpenStudy (anonymous):

since they are proportional , then you'll have this:\[\left| T \right|/ speed \] so in this case: \[343/20 = \left| 10 \right|/ \] solve for S by cross multiplying and that's it ^_^

OpenStudy (anonymous):

343/20 = 10/S *

OpenStudy (anonymous):

correct me if I'm wrong :)

OpenStudy (anonymous):

the result is 324.966 m/s

OpenStudy (anonymous):

i am thank you, leofy am very very fankful, sstarica is very good know alot than leofy, i thank you

OpenStudy (anonymous):

soory for me enlish, i am from iran

OpenStudy (anonymous):

most welcome ^_^, no it's okay, as long as we understand you then your english is perfect!

OpenStudy (anonymous):

i father is has money, so he is working a chemist for a organisation, he teach me how to make boom, and like i like it

OpenStudy (anonymous):

that's awesome, try not to create a nuclear bomb ^^" :D

OpenStudy (anonymous):

he tels me to learn this english maths and these english books, he say it is good in future

OpenStudy (anonymous):

no, i make asltrolite G

OpenStudy (anonymous):

Astrolite G

OpenStudy (anonymous):

yes, he's right :) good luck leofy!

OpenStudy (anonymous):

thankyou, i am very happy that you help me iraq people, because i know amerikans dont like iraq people, my teacher told me in sckool

OpenStudy (anonymous):

let s be the speed of light, and T the absolute temperature, then we have: s=k(sqrt(T)), where k is some constant we can find k: k=s/(sqrt(T)) => k=343/(20+273)=1.17 hence, the speed of air at -10°C is s=1.17(-10+273)=307.7m/s hopefully I am right :)

OpenStudy (anonymous):

oh there is a mistake :( k=343/sqrt(20+273)=20.04 therefor s=20.04*sqrt(-10+273)=324.97m/s :)

OpenStudy (anonymous):

hmmm, why complicate it? T and S are both proportional to each other :)

OpenStudy (anonymous):

it doesn't matter what is k because it is constant and after simplifying it get omited.so let v' be new speed.v'/v=sqrt(T'/T) where T' is new temprture in kelvin. so the last result for v' is 324.966 m/s

OpenStudy (anonymous):

it feels much better when it's more complicated, don't you agree? :)

OpenStudy (anonymous):

guys guys, they are proportional why don't you just write down the following relationship: T1/ S1 = T2/S2 simple as that ^_^

OpenStudy (anonymous):

if you can't explain it in a simpler way , then you yourself don't understand it- Albert Einstein ^_^

OpenStudy (anonymous):

wow this is a good one for Albert Einstein, although I don't agree totally with him :P

OpenStudy (anonymous):

you should, the simpler you explain it , the clearer you see the problem and the clearer the other person sees it too + the more organized. To solve a problem, we take each step at a time = this is called organization = simplicity :)

OpenStudy (anonymous):

haha ok ok I agree

OpenStudy (anonymous):

:) how great it is to find a simple solution for a complicated problem. This is life ^_^

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I think there is one little mistake in the formula you gave, that you consider S to be proportional to T, while it's proportional to its square. right?

OpenStudy (anonymous):

square root*

OpenStudy (anonymous):

true! I forgot abt it ^_^ so it'll be :\[\sqrt{T1}/\S1 =\sqrt{T2}/\S2\] ofcourse the absolute of both temperatures :)

OpenStudy (anonymous):

sqrt|T1|/S1 = sqrt|T2|/S2 ^_^

OpenStudy (anonymous):

thank you anwar :)

OpenStudy (anonymous):

today is hapy day for my, my father tell my i can go to my one training for fightrs

OpenStudy (anonymous):

yay, have fun! ^_^

OpenStudy (anonymous):

loll.. today is my worst day.. just saying not like it would matter :) anyways, good for you leofy, go beat the crap out of them :P

OpenStudy (anonymous):

leofy, you are from iraq and specifyed something about your father teaching you to make bombs, and in particular, ASTROLITE G!!! I have an idea that you ARE terrorist (at least your father is)

OpenStudy (anonymous):

we dont need terrorists here

OpenStudy (anonymous):

lol.. you are so logical :P

OpenStudy (anonymous):

oh come on, he's doing that for the love of science angoo :) no need to call him that, that's mean

OpenStudy (anonymous):

you know it's sarcasm right? :P

OpenStudy (anonymous):

i dunno, is it? .-.

OpenStudy (anonymous):

no, i googled up astrolite G, and my dear friend, that is NOT a science type of boom! see for your self

OpenStudy (anonymous):

lol my reply was

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Astrolite

OpenStudy (anonymous):

well maybe it's used as some sort of weaponary tool for his country?

OpenStudy (anonymous):

damn, half a glass of this stuff is enough to blow up a tank!!!

OpenStudy (anonymous):

AND,millitary does not use it, because it is too dangerous

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Astrolite see for yourself! I am telling you this is a terrorist kid!!!

OpenStudy (anonymous):

c'mon, I would not even bother myself to check.. it's just your thoughts being effected and controlled by the media

OpenStudy (anonymous):

see, that's the beauty of science in the field of chemistry ^_^

OpenStudy (anonymous):

maybe they are just discovering new things or trying to create new things, you never know ?

OpenStudy (anonymous):

ok, whatever, but if he learns maths and uses it to build bombs that will be used by suicide terrorists... I wash my hands

OpenStudy (anonymous):

you're saying the beauty of science in everything sstarica.. and this is the beauty of being a scientist :)

OpenStudy (anonymous):

oh come on, don't be s negatively charged angoo! I must feed you a dozen of protons to get you positive . He's just learning and he sounds like a nice guy. Maybe he's just trying out different things

OpenStudy (anonymous):

so*

OpenStudy (anonymous):

ok you wash your hands.. we will not hold you any responsibility about it. but make sure not to help him solve any problem

OpenStudy (anonymous):

lol, come on. Have faith in ppl ^_^

OpenStudy (anonymous):

how can you possibly trust yourselves without trusting one another?

OpenStudy (anonymous):

hmm....ok, ok, sorry for being too harsh, i am sorry Leofy if you are reading this

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

are you serious "BecomeMyFan" btw? because if you're, I will never be a fan of yours. (angry face)

OpenStudy (anonymous):

anwar >_>

OpenStudy (anonymous):

i was just sarcastic and messing around =D LOL

OpenStudy (anonymous):

LOL just kidding guys.. this argument really helped me change the mood I was in.

OpenStudy (anonymous):

you sounded serious angoo ^^"....and you've made me open my big mouth

OpenStudy (anonymous):

._. *dumbfounded*

OpenStudy (anonymous):

^^"

OpenStudy (anonymous):

you deserve an oscar then :P

OpenStudy (anonymous):

ahahaha... ^^" seriously.

OpenStudy (anonymous):

well, i am practising, because i want to be good at social hacking, so i gotta practice on ppl to manipulate their emotions and make them believe what i want them believe :)

OpenStudy (anonymous):

you haven't manipulated MY emotions >_>

OpenStudy (anonymous):

lol.. you know that you failed, right? :P

OpenStudy (anonymous):

yes, i am just starting out

OpenStudy (anonymous):

._. good luck?

OpenStudy (anonymous):

oh, and btw, Leofy is a fictionary character that I created

OpenStudy (anonymous):

hahah I am good at it. I can help you get better

OpenStudy (anonymous):

w...what >_>

OpenStudy (anonymous):

wes, you heard me right

OpenStudy (anonymous):

I feel like jumping out of a window right now

OpenStudy (anonymous):

we give you a face (you will never know what that means).

OpenStudy (anonymous):

gave*

OpenStudy (anonymous):

we did?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

._. ...

OpenStudy (anonymous):

didn't you?

OpenStudy (anonymous):

^^"...don't remember

OpenStudy (anonymous):

so, was i good at being Leofy, did he sound sincear, real?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

no

OpenStudy (anonymous):

he sounded so innocent >_<

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

=D

OpenStudy (anonymous):

anyways, gotta finish chapter 4 in data communications adios~ ^_^

OpenStudy (anonymous):

guys, lets start a social engineering study group

OpenStudy (anonymous):

he is so kind.. I believe he is reading what we are writing right now and he's like "what the hell is going on? :O"

OpenStudy (anonymous):

good luck sstarica :)